OpenStudy (anonymous):

solve using the substitution method 4x=y=20 2x+9y=10

OpenStudy (anonymous):

Is that tiped right?

OpenStudy (anonymous):

Because this system has no solutions

OpenStudy (anonymous):

y=20-4x 2x+9(20-4x)=10 2x+180-36x=10 -34x=-170 170=34x x=5 y=0